3. (5 pts.) Solve for χ and keep 2 decimal places x(1 +5%) = $7,000 Ans.: 4. (5pts.) Solve for χ and keep 2 decimal places = $10,000 Ans.: 5. (5 pts.) Solve for r, knowing P-360, Di-32.1 and g = 7% D1 r-g Ans.:

thank you
Answer #1

1) X (1+5%)3= $7000

X (1+ 0.05)3 = 7000

X (1.05)3 = 7000

X (1.05 x 1.05 x 1.05) = 7000

X (1.157625) = 7000

X= 7000/1.157625

X= $ 6046.86

2. X/ (1+2%)5= $10,000

X/ (1+0.02)5 = 10000

X/ (1.02)5= 10000

X/ (1.02 x 1.02 x 1.02 x 1.02 x 1.02) = 10000

X (1.10408) = 10000

X= 10000 x 1.10408

X=$ 11040.81

3. P= D1/ R-G

60 = 2.1/ (R- 7%)

60/2.1 = 1 (R-0.07)

28.5714 = 1/(R-0.07)

R - 0.07 = 1/28.5714

R - 0.07= 0.035

R = 0.035+0,07

R = 0.105 or 10.50 %